Hi,
I could use help with this quadratic equation:
c) 2x^2 + 7 = 0.75x^2 + 35
Thanks,
Louis

Answer by checkley71(8403)

2X~2+7=.75X+35 OR 1.25X~2=35-7 OR 1.25X~2=28 OR X~2=28/1.25 OR X~2=22.4 PROOF
2*22.4+7=.75*22.4+35 or 44.8+7=16.8+35 or 51.8=51.8
IF YOU HAVE TO HAVE THE VALUE OF X THE X=SQRT22.4 OR X=4.7328638